017 (part 1 of 3) 10.0 points A thin lens has a crescent shape, which i defined by two spherical surfaces A and B The radius of the surface A is RAI = 3a, and of surface B is RBI = a. The lens is submerged in a medium CB What is the magnitude of the focal length Ifl in the small angle approximation? Assume the index of refraction of the lens is 2.0 and o the medium is 1.0. 2 4 4 3 10. |f| = 4a
Answer #1

1/f = (n - 1)[1 / RA - 1 / RB]

n = 2

RA = - a

RB = 3a


1/f =(2 -1)[ - 1/a - 1/3a]

f = 3 a / 4

Ans(5)